If 4(3x+2)-(x+5)=-3 then x = ?
a) 11/6
b) -11/6
c) 6/11
d) -6/11
I'm not really understanding how to do this, I'm getting very different answers.. I think the variables are confusing me...

Answer by stanbon(75887)

If 4(3x+2)-(x+5)=-3 then x = ?
a) 11/6
b) -11/6
c) 6/11
d) -6/11
-----------------
4(3x+2)-(x+5)=-3
12x+8-x-5 = -3
11x + 3 = -3
11x = -6
x = -6/11
===========
cheers,
Stan H.